# Quality and Quantity of Machine Translation References for Automatic Metrics

### 07 Mar 2024 v4

@@ Abstract
| Vilém Zouhar, Ondřej Bojar |
| ETH Zürich, Charles University |
| vzouhar@ethz.ch, bojar@ufal.cuni.cz |
@@ 2. Related Work
The BLEU metric (Papineni et al., 2002) was intended to be used with multiple metrics, which was only rarely put in practice over the years. Nevertheless, newer and more sophisticated methods exist to incorporate them (Qin and Specia, 2015). Our results from Figure 1 confirm the older observations of Finch et al. (2004) or Bojar et al. (2013, Section 5) who study the effect of the reference count on metric performance. Finally, multiple references can be used in training better machine translation systems (Madnani et al., 2008; Zheng et al., 2018; Khayrallah et al., 2020; Mi et al., 2020) or for analyzing model uncertainty (Ott et al., 2018) or evaluation uncertainty (Zhang and Vogel, 2004, 2010; Fomicheva et al., 2020).
@@ 3. Setup
Optimal Reference Translations (ORT).
Zouhar et al. (2023) re-annotate a subset of the EnglishCzech testset from the News domain used in the WMT2020 campaign (Barrault et al., 2020). New references were created by translating the original source in 4 different human settings ranging from generic translation vendors to academics from the field of translation studies following a novel protocol leading to so-called “optimal reference translations” (ORT; Kloudová et al., 2023). This phase was followed by a human annotation and post-editing phase performed by 11 annotators of varying professionalities.
Zouhar et al. (2023) study whether the human quality of the references is really the highest achievable one. They stop short of evaluating the impact of this on machine translation evaluation. We re-purpose their data and system submissions from Barrault et al. (2020). We refer to the references, from lowest to highest quality of the source, as R1, R2, R3, and R4. Specifically, R1 to R2 come from standard translation vendors, R3 is high-quality translation vendor, and R4 is the work of translatologists (the optimal reference). See Table 1 for basic statistics.
@@
222 BLEU|#:1|c:mixed|e:yes|tok:13a|s:expchrF|#:1|c:mixed|e:yes|nc:6|nw:0|s:no
TER|#:1|c:lc|t:tercom|nr:no|pn:yes|as:no
For the metrics, we use BLEU (Papineni et al., 2002), chrF (Popović, 2015), TER (Snover et al., 2006), COMET (Rei et al., 2020), its referenceless version COMET, and its updated iteration COMET (Rei et al., 2022), and BLEURT (Sellam et al., 2020). We select these as a representative set of widely-used string-matching and trainable metrics.
@@
We focus on and evaluate metric success at the segment-level (“sentence”-level) by correlating the metric scores with human scores using Kendall’s .333 ; read more on the definition in Macháček and Bojar (2014). Each translation receives a human score and automatic metric scores which are correlated. This is the standard segment-level evaluation adopted by the WMT Metrics Shared Task (Freitag et al., 2021b, 2022, 2023). In our case (WMT2020), the human segment-level judgments were created from Direct Assessment (Graham et al., 2016) judgements following the “daRR” conversion as described by Mathur et al. (2020): Candidate translations from MT systems were scored on their own, independently of other candidates. For each pair of judgements of candidates translating the same source, we construct one golden-truth item of pairwise comparison if the two individual scores differ by more than 25% absolute. As Mathur et al. (2020), we believe that this difference is the judgement is big enough to trust the simulated pairwise comparison.
@@ 4.2. Multiple References are Useful
Biggest advantage is gained from at least three references and taking their segment-level average (max aggregation not shown).To systematically study the effect of reference count on metric performance, in Figure 1 we randomly sample references from the whole pool of original and post-edited translations, irrespective of their quality. The biggest gains in metric performance are achieved until seven references and further gains are negligible, which is in line with the observations of Bojar et al. (2013, Section 5).
@@ 4.3. References and Metric Scores
To understand the effect of different metrics, we show the average raw scores of each metric in Table 5. While it appears that the higher the raw score, the better the metric performance (low score of R1 and high scores of R3 and R4), this trend does not explain the improvements of using the post-edited versions, e.g. as R1 over R1, or R3 over R3. In fact, the post-edited versions always lead to lower raw scores. This could be the result of either further translations shifts as the post-edits are based on a translation and not the source or additional (fully justified) corrections in the references which lead to fewer matches with the candidates.
@@ 4.4. Allocating a Budget for References
Usually, it is simple to gather many source sentences and let multiple systems translate them. Evaluating all of them using human annotators is unattainable but running automatic metrics is not. However, these require references, which are also costly. It remains unclear how many references and of what quality level should we obtain for a given budget to achieve the most reliable automatic quality assessment.
Mixing references does not hurt any metric.Input: Source segments , levels , cost function , utility function
@@
To assess what types of configurations of references can lead to the most reliable automatic evaluation, we first validate if references can be meaningfully be mixed. For example, if it is viable that 75% of the sources can have references from a cheaper vendor R1 and 25% from a higher-quality but more expensive vendor R3. This is different from Table 4 where each segment had exactly two references from the same two sources. In Section 4.3, we show that using lower-quality references R1 leads to lower absolute metric scores (e.g. BLEU = 24.2) as opposed to higher-quality ones R3 (e.g. BLEU = 37.1). This holds across all metrics. Bojar et al. (2010) observe that lower BLEU scores are less reliable, but they refer to the range of BLEU 20. It is thus questionable if BLEUs at 20–40 correlate differently with human MT quality judgements. In Figure 2, we mix some of the references together for evaluation, but staying at single-reference evaluation. Despite the varying absolute scores of metrics under different references, as explored in Section 4.3, mixing of multiple references leads to an almost perfectly linear combination of the endpoint metric performances. The biggest gains in this respect are obtained by BLEURT, chrF and BLEU, while COMET is almost unaffected. There is no formal guarantee that the mix of score distributions will not lower the overall Kendall’s . Nevertheless, a positive conclusion is that if there is budget to only translate 25% of segments with high quality, it should be done and it can only improve the overall evaluation reliability.
@@
We provide a heuristic-based stochastic algorithm to find an assignment of source segments to be translated by vendors of different costs and qualities within a specific budget. For the current dataset, we set the cost of a segment in R1, R2, R3, and R4 to be 1, 1, 2, and 3, respectively. Their quality (or “fitness” for the purpose of automatic evaluation) were set to 1, 2, 4, and 3 based on our observations in Section 4.1. Algorithm 1 contains a hyperparameter that controls whether the budget will be allocated more towards having multiple references per-segment or more towards having fewer but higher-quality references per-segment, and the temperature than controls the sampling.
@@
With a limited budget, e.g. or , it makes more sense to add some references of a higher quality rather than covering the whole test set with a second reference.
With more budget available, multiple references per segment become more beneficial.
@@ 6. Conclusion
Figure 5: Average number of references per one segment allocated by Algorithm 1 with (top) and (bottom).
@@
| 10 | 78 | K evropské špičce nově patří i tři skotští studenti |
| 23 | 120 | Tři skotští studenti se umístili mezi nejlepšími v Evropě |
| 23 | 121 | Tři skotští studenti mezi nejlepšími v Evropě |
| 28 | 116 | Tři skotští studenti byli oceněni jako jedni z nejlepších v Evropě |
| 28 | 115 | Tři skotští studenti byli jmenováni jako jedni z nejlepších v Evropě |
| 28 | 114 | Tři skotští studenti byli vyhlášeni jako jedni z nejlepších v Evropě |
| 32 | 117 | Tři skotští studenti byli jmenováni jednimi z nejlepších v Evropě |
| 37 | 122 | Tři skotští studenti byli jmenováni mezi nejlepšími v Evropě |
| 43 | 125 | Tři skotští studenti se zařadili mezi nejlepší v Evropě |
| 43 | 121 | Tři skotští studenti patří mezi nejlepší v Evropě. |
| 43 | 122 | Tři skotští studenti patří mezi nejlepší v Evropě |
| 60 | 127 | Tři skotští studenti zařazeni mezi nejlepší v Evropě |
| 100 | 131 | Tři skotští studenti byli zařazeni mezi nejlepší v Evropě |
@@
| 4 | -42 | Sony a Disney točí třetí film o Spidermanovi |
| 4 | -33 | Sony a Disney točí třetí film o Spider-Manovi |
| 8 | -9 | Sony a Disney pracují na třetím filmu o Spider-Manovi |
| 8 | -4 | Sony a Disney pokračují v práci na třetím filmu o Spider-Manovi |
| 8 | 1 | Sony a Disney opět pracují na třetím filmu o Spider-Manovi |
| 8 | 15 | Sony a Disney spolupracují na třetím filmu o Spider-Manovi |
| 4 | 16 | Sony a Disney budou spolupracovat při natáčení třetího filmo o Spider-manovi |
| 8 | 28 | Sony a Disney opět spolupracují na třetím filmu o Spider-Manovi |
| 8 | 30 | Sony a Disney budou spolupracovat na třetím filmu o Spider-Manovi |
| 8 | 35 | Sony a Disney budou opět spolupracovat na třetím filmu o Spider-Manovi |
| 17 | 52 | Disney bude znovu spolupracovat se společností Sony na třetím filmu Spider-Man |
| 10 | 64 | Disney bude se Sony dál pracovat na třetím filmu se Spider-Manem |
| 50 | 73 | Disney bude spolupracovat se Sony na třetím sólovém filmu o Spider-Manovi |
| 75 | 99 | Disney se vrací, bude spolupracovat se Sony na třetím filmu o Spider-Manovi |
| 78 | 106 | Disney se vrací, bude spolupracovat se Sony na třetím sólovém filu Spider-Man. |
| 79 | 108 | Disney se vrací, bude spolupracovat se Sony na třetím Spider-Man filmu |
| 100 | 121 | Disney se vrací, bude spolupracovat se Sony na třetím sólovém Spider-Man filmu |
@@ Acknowledgements
- Fomicheva et al. (2020) Marina Fomicheva, Lucia Specia, and Francisco Guzmán. 2020. Multi-hypothesis machine translation evaluation. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics.
@@
- Papineni et al. (2002) Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002. Bleu: a method for automatic evaluation of machine translation. In Proceedings of the 40th Annual Meeting of the Association for Computational Linguistics, 311–318. Association for Computational Linguistics.
@@
- Qin and Specia (2015) Ying Qin and Lucia Specia. 2015. Truly exploring multiple references for machine translation evaluation. In Proceedings of the 18th Annual Conference of the European Association for Machine Translation. European Association for Machine Translation.
@@
- Sellam et al. (2020) Thibault Sellam, Dipanjan Das, and Ankur Parikh. 2020. BLEURT: Learning robust metrics for text generation. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, 7881–7892. Association for Computational Linguistics.

### 10 Apr 2024 v5

@@ Abstract
Vilém Zouhar Ondřej Bojar |
ETH Zürich Charles University |
vzouhar@ethz.ch bojar@ufal.cuni.cz |
@@ 2. Related Work
The BLEU metric (Papineni et al., 2002) was intended to be used with multiple metrics, which was only rarely put in practice over the years. Nevertheless, newer and more sophisticated methods exist to incorporate them (Qin and Specia, 2015). Our results from Figure 1 confirm the older observations of Finch et al. (2004) or Bojar et al. (2013, Section 5) who study the effect of the reference count on metric performance. Finally, multiple references can be used in training better machine translation systems (Madnani et al., 2008; Zheng et al., 2018; Khayrallah et al., 2020; Mi et al., 2020) or for analyzing model uncertainty (Ott et al., 2018) or evaluation uncertainty (Zhang and Vogel, 2004, 2010; Fomicheva et al., 2020). It is also used outside of machine translation for measuring consensus (Vedantam et al., 2015).
@@ 3. Setup
Optimal Reference Translations.
Zouhar et al. (2023) re-annotate a subset of the EnglishCzech testset from the News domain of the WMT2020 campaign (Barrault et al., 2020). New references were created by translating the original source in 4 different human settings ranging from generic translation vendors to translatology academics following a novel protocol leading to so-called “optimal reference translations” (Kloudová et al., 2023). This phase was followed by a human annotation and post-editing phase performed by 11 annotators of varying professionalities.
Zouhar et al. (2023) study whether the human quality of the references is really the highest achievable one. They stop short of evaluating the impact of this on machine translation evaluation. We re-purpose their data and system submissions from Barrault et al. (2020). We refer to the references, from lowest to highest quality of the source, as R1, R2, R3, and R4. Specifically, R1 to R2 come from standard translation vendors,222Nevertheless, based on observations of Kloudová et al. (2021), R1 are to a large extent post-edits of one of the participating systems. R3 is high-quality translation vendor, and R4 is the work of translatologists (the optimal reference). See Table 1 for basic statistics.
chrF|#:1|c:mixed|e:yes|nc:6|nw:0|s:no
TER|#:1|c:lc|t:tercom|nr:no|pn:yes|as:no
@@
33footnotemark: 3For the metrics, we use BLEU (Papineni et al., 2002), chrF (Popović, 2015), TER (Snover et al., 2006), COMET (Rei et al., 2020), its referenceless version COMET, and its updated iteration COMET (Rei et al., 2022), and BLEURT (Sellam et al., 2020). We select these as a representative set of widely-used string-matching and trainable metrics.
@@
We focus on and evaluate metric success at the segment-level (“sentence”-level) by correlating the metric scores with human scores using Kendall’s .444 ; read more on the definition in Macháček and Bojar (2014). Each translation receives a human score and automatic metric scores which are correlated. This is the standard segment-level evaluation adopted by the WMT Metrics Shared Task (Freitag et al., 2021b, 2022, 2023). In our case (WMT2020), the human segment-level judgments were created from Direct Assessment (Graham et al., 2016) judgements following the “daRR” conversion as described by Mathur et al. (2020): Candidate translations from MT systems were scored on their own, independently of other candidates. For each pair of judgements of candidates translating the same source, we construct one golden-truth item of pairwise comparison if the two individual scores differ by more than 25% absolute. As Mathur et al. (2020), we believe that this difference in the judgement is big enough to trust the simulated pairwise comparison.
@@ 4.2. Multiple References are Useful
To systematically study the effect of reference count on metric performance, in Figure 1 we randomly sample references from the whole pool of original and post-edited translations, irrespective of their quality. The biggest gains in metric performance are achieved until seven references and further gains are negligible, which is in line with the observations of Bojar et al. (2013, Section 5).
@@ 4.3. References and Metric Scores
To understand the effect of different metrics, we show the average raw scores of each metric in Table 5. While it appears that the higher the raw score, the better the metric performance (low score of R1 and high scores of R3 and R4), this trend does not explain the improvements of using the post-edited versions, e.g. as R1 over R1, or R3 over R3. In fact, the post-edited versions always lead to lower raw scores. This could be the result of either further translation shifts as the post-edits are based on a translation and not the source or additional (fully justified) corrections in the references which lead to fewer matches with the candidates.
@@ 4.4. Allocating a Budget for References
Usually, it is simple to gather many source sentences and let multiple systems translate them. Evaluating all of them using human annotators is unattainable but running automatic metrics is not. However, these require references, which are also costly. It remains unclear how many references and of which quality to obtain to achieve the most reliable automatic quality assessment under a given budget.
Input: Source segments , levels , cost function , utility function
@@
To assess what types of configurations of references can lead to the most reliable automatic evaluation, we first validate if references can be meaningfully mixed. For example, if it is viable that 75% of the sources can have references from a cheaper vendor R1 and 25% from a higher-quality but more expensive vendor R3. This is different from Table 4 where each segment had exactly two references from the same two sources. In Section 4.3, we show that using lower-quality references R1 leads to lower absolute metric scores (e.g. BLEU = 24.2) as opposed to higher-quality ones R3 (e.g. BLEU = 37.1). This holds across all metrics. Bojar et al. (2010) observe that lower BLEU scores are less reliable, but they refer to the range of BLEU 20. It is thus questionable if BLEUs at 20–40 correlate differently with human MT quality judgements. In Figure 2, we mix some of the references together for evaluation, but staying at single-reference evaluation. Despite the varying absolute scores of metrics under different references, as explored in Section 4.3, mixing of multiple references leads to an almost perfectly linear combination of the endpoint metric performances. The biggest gains in this respect are obtained by BLEURT, chrF and BLEU, while COMET is almost unaffected. There is no formal guarantee that the mix of score distributions will not lower the overall Kendall’s . Nevertheless, a positive conclusion is that if there is budget to only translate 25% of segments with high quality, it should be done and it can only improve the overall evaluation reliability.
@@
We provide a heuristic stochastic algorithm to find an assignment of source segments to be translated by vendors of different costs and qualities within a specific budget. For the current dataset, we set the cost of a segment in R1, R2, R3, and R4 to be 1, 1, 2, and 3, respectively. Their quality (or “fitness” for the purpose of automatic evaluation) were set to 1, 2, 4, and 3 based on our observations in Section 4.1. Algorithm 1 contains a hyperparameter that controls whether the budget will be allocated more towards having multiple references per-segment or more towards having fewer but higher-quality references per-segment, and the temperature than controls the the sampling randomness.
@@ 6. Conclusion
![[Uncaptioned image]](x21.png)
![[Uncaptioned image]](x23.png)
@@
10 | 78 | K evropské špičce nově patří i tři skotští studenti |
23 | 120 | Tři skotští studenti se umístili mezi nejlepšími v Evropě |
23 | 121 | Tři skotští studenti mezi nejlepšími v Evropě |
28 | 116 | Tři skotští studenti byli oceněni jako jedni z nejlepších v Evropě |
28 | 115 | Tři skotští studenti byli jmenováni jako jedni z nejlepších v Evropě |
28 | 114 | Tři skotští studenti byli vyhlášeni jako jedni z nejlepších v Evropě |
32 | 117 | Tři skotští studenti byli jmenováni jednimi z nejlepších v Evropě |
37 | 122 | Tři skotští studenti byli jmenováni mezi nejlepšími v Evropě |
43 | 125 | Tři skotští studenti se zařadili mezi nejlepší v Evropě |
43 | 121 | Tři skotští studenti patří mezi nejlepší v Evropě. |
43 | 122 | Tři skotští studenti patří mezi nejlepší v Evropě |
60 | 127 | Tři skotští studenti zařazeni mezi nejlepší v Evropě |
100 | 131 | Tři skotští studenti byli zařazeni mezi nejlepší v Evropě |
![[Uncaptioned image]](x25.png)
@@
4 | -42 | Sony a Disney točí třetí film o Spidermanovi |
4 | -33 | Sony a Disney točí třetí film o Spider-Manovi |
8 | -9 | Sony a Disney pracují na třetím filmu o Spider-Manovi |
8 | -4 | Sony a Disney pokračují v práci na třetím filmu o Spider-Manovi |
8 | 1 | Sony a Disney opět pracují na třetím filmu o Spider-Manovi |
8 | 15 | Sony a Disney spolupracují na třetím filmu o Spider-Manovi |
4 | 16 | Sony a Disney budou spolupracovat při natáčení třetího filmo o Spider-manovi |
8 | 28 | Sony a Disney opět spolupracují na třetím filmu o Spider-Manovi |
8 | 30 | Sony a Disney budou spolupracovat na třetím filmu o Spider-Manovi |
8 | 35 | Sony a Disney budou opět spolupracovat na třetím filmu o Spider-Manovi |
17 | 52 | Disney bude znovu spolupracovat se společností Sony na třetím filmu Spider-Man |
10 | 64 | Disney bude se Sony dál pracovat na třetím filmu se Spider-Manem |
50 | 73 | Disney bude spolupracovat se Sony na třetím sólovém filmu o Spider-Manovi |
75 | 99 | Disney se vrací, bude spolupracovat se Sony na třetím filmu o Spider-Manovi |
78 | 106 | Disney se vrací, bude spolupracovat se Sony na třetím sólovém filu Spider-Man. |
79 | 108 | Disney se vrací, bude spolupracovat se Sony na třetím Spider-Man filmu |
100 | 121 | Disney se vrací, bude spolupracovat se Sony na třetím sólovém Spider-Man filmu |
@@ Acknowledgements
- Fomicheva et al. (2020) Marina Fomicheva, Lucia Specia, and Francisco Guzmán. 2020. Multi-hypothesis machine translation evaluation. In Proceedings of the 58th Annual Meeting of ACL.
@@
- Kloudová et al. (2021) Věra Kloudová, Ondřej Bojar, and Martin Popel. 2021. Detecting post-edited references and their effect on human evaluation. In Proceedings of the Workshop on Human Evaluation of NLP Systems (HumEval), 114–119, Stroudsburg, USA. Association for Computational Linguistics.
@@
- Papineni et al. (2002) Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002. Bleu: a method for automatic evaluation of machine translation. In Proceedings of the 40th Annual Meeting of ACL, 311–318. Association for Computational Linguistics.
@@
- Qin and Specia (2015) Ying Qin and Lucia Specia. 2015. Truly exploring multiple references for machine translation evaluation. In Proceedings of the 18th Annual Conference of the European Association for Machine Translation.
@@
- Sellam et al. (2020) Thibault Sellam, Dipanjan Das, and Ankur Parikh. 2020. BLEURT: Learning robust metrics for text generation. In Proceedings of the 58th Annual Meeting of ACL, 7881–7892. Association for Computational Linguistics.
@@
- Vedantam et al. (2015) Ramakrishna Vedantam, C Lawrence Zitnick, and Devi Parikh. 2015. Cider: Consensus-based image description evaluation. In Proceedings of the IEEE conference on computer vision and pattern recognition, 4566–4575.